Calculate the reaction entropy (S°rxn) for the following reaction: 2 LiOH + CO2 Li2CO3 + H2O


1
Expert's answer
2022-05-11T03:16:03-0400

ΔS°rxn = ΔS(H2O) + ΔS(Li2CO3) - ΔS(CO2) - 2ΔS(LiOH)

ΔS°rxn= 69.91 + 90.17 - 213.68 - 2(42.8) = - 139.2 J/K  (decrease in entropy)